# Hierarchical Clustering With Confidence

### v1

@@ Abstract
We propose a simple randomization scheme together with a method for constructing a valid p-value at each node of the hierarchical clustering dendrogram that quantifies evidence against performing the greedy merge. Our test controls the Type I error rate, works with any hierarchical linkage without case-specific derivations, and simulations show it is substantially more powerful than existing selective inference approaches. To demonstrate the practical utility of our p-values, we develop an adaptive $\alpha$-spending procedure that estimates the number of clusters, with a probabilistic guarantee on overestimation. Experiments on simulated and real data show that this estimate yields powerful clustering and can be used, for example, to assess clustering stability across multiple runs of the randomized algorithm.
@@ 1 Introduction
Agglomerative hierarchical clustering is one of the most widely used approaches for exploring relationships among observations in a dataset. By producing nested clusters, it allows researchers to visualize how observations group together across different levels of similarity, without the need to specify the number of clusters in advance. Owing to its ease of implementation and the convenient visualization of its results, this class of clustering methods has found extensive applications in diverse fields such as genomics ((Eisen et al., 1998)), neuroscience (Moreno-Dominguez et al. (2014); Blumensath et al. (2013)), ecology ((Zolfaghari et al., 2019)), network analysis (Shen et al. (2009); Chen et al. ), and text mining ((Steinbach et al., 2000)).
@@
Figure 1: Figure 1: (a): A two-dimensional example with $n=30$ points across two true clusters; (b): a dendrogram resulting from complete-linkage hierarchical clustering of the example data; (c): a dendrogram resulting from complete-linkage randomized hierarchical clustering of the example data (using a randomization parameter $\tau=0.10$, defined later in Section 2); (d): A histogram for the estimated number of clusters $\hat{K}$ based on the proposed $\alpha$-spending procedure.
@@
Although a complete nested clustering structure can be obtained from the tree-like dendrogram, hierarchical clustering is, at its core, a greedy algorithm. Without principled approaches to quantify evidence against greedy merges or to identify the appropriate level at which to cut the dendrogram, distinguishing genuine clusters from patterns arising purely out of sampling variation becomes difficult. For instance, because of its inherently greedy approach, hierarchical clustering results can be highly sensitive to even small perturbations in the data. As a general strategy to evaluate clustering results in the absence of labels or costly validation data, several papers (von Luxburg (2010); Liu et al. (2022); Lange et al. (2004)) have advocated the concept of clustering stability. The underlying idea is that genuine clustering structure should remain stable or reproducible when applied to perturbed datasets that are quite similar to the original data. For this reason, Monti et al. (2003) and Hennig (2007) have suggested introducing randomization into the algorithm to reveal the stability with which pairs of points are clustered.
@@
In this paper, we show how randomizing a hierarchical clustering algorithm can be useful not just for measuring stability but also for constructing valid hypothesis testing procedures based on the clustering results. We propose a simple randomization scheme together with a method for forming a valid p-value at each node of the hierarchical clustering dendrogram, which quantifies the statistical evidence against performing the greedy merge. Panel (c) of Figure 1 displays a dendrogram produced by one run of our randomized complete-linkage hierarchical clustering algorithm. It matches the dendrogram in panel (b) in terms of the two real clusters in the data, while scrambling the order in which points within each cluster are merged.
@@
Our randomization approach is motivated by recent inferential advances for decision trees (Bakshi et al., 2024; Bakshi and Panigrahi, 2025), in which conventional splits are replaced with randomized, softmax-based splits to account for the adaptive tree structure during inference. The scope of our approach, however, is not limited to performing post-hoc inference on clusters obtained from a fixed level of the algorithm. While hierarchical clustering does not require a prespecified number of clusters, interpreting its output in practice requires choosing an appropriate dendrogram level or cut height to extract and assess the resulting cluster structure. To address this practical issue, we develop an adaptive $\alpha$-spending procedure based on our valid p-values to estimate the number of clusters, providing a probabilistic guarantee against overestimation. Continuing with the example in Figure 1, we apply our proposed method to estimate the number of clusters, $\widehat{K}$, over $100$ runs of our randomized algorithm; panel (d) displays the empirical histogram of $\widehat{K}$, which has a strong mode at $2$, the true number of clusters.
@@ 1.1 Related work and contributions
Determining whether two clusters identified by a clustering algorithm are genuinely distinct, when the same data are used both to discover and to validate these clusters, is an instance of double dipping on the data. Selective inferential methods address double-dipping by explicitly accounting for the dependence created when the same data are used for both discovery and validation. Novel selective inference methodologies have been proposed by Gao et al. (2024); Yun and Barber (2023) to enable valid post-hoc inference on clustering results obtained from agglomerative clustering. Beyond hierarchical clustering, Chen and Witten (2023) develop a selective inferential approach to validate clusters identified by the K-means algorithm and Chen and Gao (2025) propose a test for differences in a single feature between a pair of clusters. By conditioning on the output of the clustering algorithm, these approaches yield valid p-values at a prespecified dendrogram level, controlling the Type I error rate. Zhang et al. (2019) address a similar problem using sample splitting, which conditions on more information than necessary compared with the approaches in Gao et al. (2024) and Yun and Barber (2023). Other prior work, such as Liu et al. (2008); Maitra et al. (2012); Suzuki and Shimodaira (2006), primarily test the global null hypothesis that no clustering structure exists in the data, which differs fundamentally in spirit from the hypotheses considered in selective inference approaches.
To test the inherently data-dependent hypotheses at each merge in the dendrogram, we condition on the clustering outcome, following the conditional framework in existing selective inference methodologies for clustering. However, because we add external randomization to the clustering algorithm, our approach to generating p-values is substantially different from previous approaches and offers several advantages, as outlined below.
@@
    Our approach, as demonstrated through simulations, produces substantially more powerful p-values than the non-randomized methods when testing mean differences between a pair of clusters at a fixed dendrogram height. Intuitively, by introducing randomization into the clustering procedure, we are saving more information for the inference stage. This often comes at little cost to the clustering quality when only a small amount of randomization is introduced into the algorithm, as seen for example in Figure 1. Gains in inferential power due to external randomization has been well-documented in supervised learning contexts. Take, for example, data-carving Fithian et al. (2014); Panigrahi (2023), which selects models using a random subsample like sample-splitting; but, unlike the sample-splitting approach, data-carving uses the full data for inference. Approaches using other types of randomization schemes are described in Tian and Taylor (2018); Panigrahi and Taylor (2023); Panigrahi et al. (2024); Kivaranovic and Leeb (2024); Huang et al. (2025), among others.
@@
    Furthermore, unlike randomized methods based on data-splitting, like the UV decomposition, fission, or thinning (Rasines and Young, 2022; Leiner et al., 2025; Dharamshi et al., 2025), that base inference exclusively on holdout data, our approach reuses the data employed in clustering. In particular, with a small amount of randomization, our approach produces a slight perturbation of the clustering results on the original data while still using the full dataset for testing.
@@
    The p-values obtained from our randomized approach can be used in conjunction with any hierarchical clustering linkage to produce valid inference, without the need for: (i) case-by-case derivations of the conditioning set or (ii) potentially expensive Monte Carlo or resampling-based approximations to compute the p-values. For example, in complete-linkage hierarchical clustering (which was used in Figure 1), the tests in Gao et al. (2024); Yun and Barber (2023) rely on Monte Carlo approximations to compute p-values, whereas our method does not resort to such sampling-based approximations.
@@
    Adding to the flexibility of our approach, the proposed randomization scheme can be readily applied to obtain adjusted distributions of test statistics for different types of tests, while accounting for the data-dependent nature of clustering outputs. Examples include tests based on the adjusted chi-squared test statistic in Gao et al. (2024) and the adjusted F test statistic in Yun and Barber (2023).
@@
  • In Section 2, we introduce the randomized hierarchical clustering algorithm (Algorithm 1), which performs clustering using a dissimilarity matrix that is updated at each merge, as in standard methods.
@@
  • Simulation results in Section 5 demonstrate that: (i) our randomized approach yields valid p-values at each merge in the dendrogram, (ii) tests based on these p-values achieve higher power than existing selective inference methods at a fixed dendrogram height, while not sacrificing clustering performance when only a small amount of randomization is used, and (iii) the proposed $\alpha$-spending procedure to estimate the dendrogram cut height yields a powerful clustering method that outperforms the standard clustering approach paired with the popular gap statistic (Tibshirani et al., 2001). In a real data analysis, we illustrate how the number of clusters estimated using our method can be employed to assess the stability of clustering results across multiple runs of our randomized Algorithm 1.
@@ 2.1 Proposed randomization scheme
We begin by introducing some notation. Let $X\in\mathbb{R}^{n\times p}$ denote the random data matrix with $n$ observations and $p$ features, and let $X_{o}$ denote its realized value; i.e., we observe $X=X_{o}$. For our hierarchical clustering algorithm, at step $t$, let $\mathcal{C}^{(t)}$ denote the set of all existing clusters, and let $\mathcal{M}^{(t)}$ denote the set of all candidate merges, where each candidate merge $M=(C_{1},C_{2})\in\mathcal{M}^{(t)}$ consists of a pair of distinct clusters $(C_{1},C_{2})$ from the set $\mathcal{C}^{(t)}$. We denote by $M^{(t)}\in\mathcal{M}^{(t)}$ the merge selected by the randomized agglomerative clustering algorithm at step $t$, which depends on both the random data matrix and the external randomization. These merges also depend on the dissimilarity between any pair of clusters $M=(C_{1},C_{2})$, which we denote by $d(M;X)$; the specific form of $d(M;\cdot)$ is determined by the linkage function employed in the hierarchical clustering algorithm. Finally, let $D^{(t)}(X)=\{d(M;X):M\in\mathcal{M}^{(t)}\}$ be the set of all dissimilarities between every possible pair of clusters or merges considered at step $t$.
To perform randomized hierarchical clustering, we propose the following randomization scheme. Given $X=X_{o}$, the set of dissimilarities at all possible merges, we sample the greedy merge $M^{(t)}=\left({C}^{(t)}_{1},{C}^{(t)}_{2}\right)$ at step $t$ from a discrete probability distribution, instead of deterministically selecting the pair of clusters from $\mathcal{C}^{(t)}$ with the smallest dissimilarity. In particular, the randomized merge variable $M^{(t)}$ assumes values in the set of all possible merges, $\mathcal{M}^{(t)}$, with probabilities given by
$\mathbb{P}_{X_{o}}^{(t)}=\left\{p^{(t)}(M;X_{o}):M\in\mathcal{M}^{(t)}\right\}$
where
$p^{(t)}(M;X_{o}):=\mathbb{P}(M^{(t)}=M|X=X_{o})=\dfrac{\exp\left(-\dfrac{1}{\tau^{(t)}}d(M;X_{o})\right)}{\displaystyle\sum_{M^{\prime}\in\mathcal{M}^{(t)}}\exp\left(-\dfrac{1}{\tau^{(t)}}d(M^{\prime};X_{o})\right)},$ (1)
and $\tau^{(t)}>0$ controls the amount of randomization applied at step $t$ of the algorithm. More succinctly, the probability vector $\mathbb{P}_{X_{o}}^{(t)}$ is the softmax of $D^{(t)}(X)/\tau^{(t)}$.
By sampling a realization of $M^{(t)}$ from the discrete distribution in (1), the algorithm at step $t$ performs a randomized version of the non-randomized or deterministic cluster-merging process. This process is repeated at each step $t$ of the algorithm, where independent sampling is performed at each merge conditioned on the data, continuing until all clusters are merged into a single cluster or a predefined number of clusters is reached.
The randomization scheme introduced above closely resembles the exponential mechanism by McSherry and Talwar (2007), which is used in differential privacy to randomly sample an item from a discrete set while providing privacy guarantees. Recent work by Zhang et al. (2024) uses a similar type of sampling probabilities as weights in a weighted average-type procedure to obtain stability, which in turn is used to establish the asymptotic normality of cross-validated risks. A key distinction of our randomized approach from this work and from the broader literature on differential privacy and algorithmic stability is that we neither require nor seek to achieve any form of privacy or stability in order to construct valid testing procedures, as described later in Section 3.
@@ 2.2 Amount of randomization
At each step of the bottom-up algorithm, one must determine the amount of randomization required to select the winning merge. For $D^{(t)}(X)=D^{(t)}(X_{o})$, the amount of randomization at each iteration, $\tau^{(t)}=\tau^{(t)}(D^{(t)}(X_{o}))$, is parameterized by a single scalar $\tau$, fixed as
$\tau^{(t)}:=\tau\cdot\frac{1}{|\mathcal{M}^{(t)}|}\sum_{M^{\prime}\in\mathcal{M}^{(t)}}d(M^{\prime};X_{o}).$ (2)
This data-driven specification of $\tau^{(t)}$, defined as a scalar multiple of the average distance among possible merges at step $t$, ensures that the clustering structure produced by our randomized algorithm remains invariant under rescaling of the data, a property also exhibited by the deterministic version.
Before proceeding, we make the following observation about our randomization scheme, obtained by substituting (2) into the probability distribution defined in (1). By definition, the random merge variable $M^{(t)}$ depends on the data matrix exclusively through the collection of all possible dissimilarities $D^{(t)}(X)$ at that step. This observation is formalized in Lemma 1.
Lemma 1.
The merge variable $M^{(t)}$ is conditionally independent of $X$ given $D^{(t)}(X)=D^{(t)}(X_{o})$, i.e.,
$\mathbb{P}(M^{(t)}=M|D^{(t)}(X)=D^{(t)}(X_{o}),X=X_{o})=\mathbb{P}(M^{(t)}=M|D^{(t)}(X)=D^{(t)}(X_{o})).$
The proof for this result is included in Appendix B.1.
The smaller the value of $\tau$, the lower the amount of randomization in our hierarchical clustering algorithm. Our next result, Proposition 1, establishes that as $\tau\to 0$, the greedy merge performed at step $t$ by our randomized clustering algorithm coincides, with probability approaching $1$, with the merge produced by its deterministic analog using the same linkage criterion. Corroborating this result, our numerical experiments in Section 5.1 show that small values of $\tau$ lead to clustering results that closely resemble the clustering output on the full data in terms of standard clustering evaluation metrics.
@@
Denote by
$\mathcal{M}^{*}_{o}=\displaystyle\operatorname*{arg\,min}_{M\in\mathcal{M}^{(t)}}d(M;X_{o})$
the set of merges with the smallest dissimilarity, as identified by the deterministic hierarchical clustering algorithm at step $t$. Then, we have that
$\lim_{\tau\rightarrow 0}\mathbb{P}(M^{(t)}\in\mathcal{M}^{*}_{o}|X=X_{o})=1.$
A detailed proof for this result is deferred to Appendix B.2.
Algorithm 1 summarizes the proposed randomized agglomerative clustering procedure, which produces $K$ clusters by performing nKn-K merges. In the special case where $K=1$, the algorithm proceeds until all observations are merged into a single cluster.
@@
1:Input: Pairwise dissimilarity matrix $D(X_{o})$, number of clusters $K$, linkage function $d(\cdot;X_{o})$, parameter controlling the amount of randomization $\tau$
2:Initialization: Each data point as its own singleton cluster:
$\mathcal{C}^{(0)}=\left\{\{1\},\{2\},\dots,\{n\}\right\}$
@@
4:  For each pair of cluster $M\in\mathcal{M}^{(t)}$, compute $d(M;X_{o})$ using $D(X_{o})$ and $d(\cdot;X_{o})$
5:  Set $\tau^{(t)}:=\tau\cdot\dfrac{1}{|\mathcal{M}^{(t)}|}\sum_{M\in\mathcal{M}^{(t)}}d(M;X_{o})$
6:  Calculate $p^{(t)}(M;X_{o})$, as defined in (1), for all $M\in\mathcal{M}^{(t)}$
7:  Sample the winning merge $M_{o}^{(t)}=\left(C^{(t)}_{1,o},C^{(t)}_{2,o}\right)$ from
$\mathbb{P}_{X_{o}}^{(t)}=\left\{p^{(t)}(M;X_{o}):M\in\mathcal{M}^{(t)}\right\}$
8:  Merge the winning pair of clusters and update the collection of clusters:
$\mathcal{C}^{(t+1)}\leftarrow\mathcal{C}^{(t)}\setminus\{C^{(t)}_{1,o},C^{(t)}_{2,o}\}\cup\left\{C^{(t)}_{1,o}\cup C^{(t)}_{2,o}\right\}$
9:end for
10:Output: Cluster labels for each data point

### v2

@@ Abstract
We propose a simple randomization scheme together with a method for constructing a valid p-value at each node of the hierarchical clustering dendrogram that quantifies evidence against performing the merge. Our test controls the Type I error rate, works with any hierarchical linkage without case-specific derivations, and simulations show it is substantially more powerful than existing selective inference approaches. To demonstrate the practical utility of our p-values, we develop an adaptive $\alpha$-spending procedure that estimates the number of clusters, with a probabilistic guarantee on overestimation. Experiments on simulated and real data show that this estimate yields powerful clustering and can be used, for example, to assess clustering stability across multiple runs of the randomized algorithm.
Keywords: Exponential mechanism, Hypothesis testing, Randomization, Selective inference, Stability, Type I error
@@ 1 Introduction
Agglomerative hierarchical clustering is one of the most widely used approaches for exploring relationships among observations in a dataset. By producing nested clusters, it allows researchers to visualize how observations group together across different levels of similarity, without specifying the number of clusters in advance. Owing to its ease of implementation and the convenient visualization of its results, this class of clustering methods has found extensive applications in diverse fields such as genomics ([Eisen et al., 1998]), neuroscience ([Moreno-Dominguez et al., 2014]), ecology ([Zolfaghari et al., 2019]), network analysis (Shen et al. [2009], Chen et al. ), and text mining ([Steinbach et al., 2000]).
@@
Figure 1: (a): A two-dimensional example with $n=30$ points across two true clusters; (b): a dendrogram resulting from complete-linkage hierarchical clustering of the example data; (c): a dendrogram resulting from complete-linkage randomized hierarchical clustering of the example data (using a randomization parameter $\tau^{*}=0.10$, defined later in Section 2); (d): A histogram for the estimated number of clusters $\hat{K}$ based on the proposed $\alpha$-spending procedure.
@@
Although a complete nested clustering structure can be obtained from the tree-like dendrogram, hierarchical clustering is, at its core, a greedy algorithm. Without principled approaches to quantify evidence against greedy merges or to identify the appropriate cut level, it becomes difficult to distinguish genuine clusters from patterns arising purely out of sampling variation. For instance, hierarchical clustering results can be highly sensitive to small perturbations in the data due to its greedy nature. To evaluate clustering results in the absence of labels or costly validation data, several papers (von Luxburg [2010], Liu et al. [2022], Lange et al. [2004]) have advocated the concept of clustering stability. The underlying idea is that genuine clustering structure should remain reproducible under small perturbations. To this end, Monti et al. [2003] and Hennig [2007] have suggested introducing randomization into the algorithm to reveal the stability of pairwise cluster assignments.
@@
In this paper, we show how randomizing a hierarchical clustering algorithm can be useful not just for measuring stability but also for constructing valid hypothesis testing procedures based on the clustering results. We propose a simple randomization scheme together with a method for forming a valid p-value at each node of the dendrogram, which quantifies the statistical evidence against performing the merge. Panel (c) of Figure 1 displays a dendrogram produced by one run of our randomized complete-linkage hierarchical clustering algorithm. It matches the dendrogram in panel (b) in terms of the two real clusters in the data, while scrambling the order in which points within each cluster are merged.
@@
Our randomization approach is motivated by recent inferential advances for decision trees [Bakshi et al., 2024, Bakshi and Panigrahi, 2025], in which conventional splits are replaced with randomized, softmax-based splits to account for the adaptive tree structure during inference. The scope of our approach, however, is not limited to performing post-hoc inference on clusters obtained from a fixed level of the algorithm. While hierarchical clustering does not require a prespecified number of clusters, interpreting its output in practice requires choosing an appropriate dendrogram level or cut height to extract and assess the resulting cluster structure. To address this practical issue, we develop an adaptive $\alpha$-spending procedure based on our valid p-values to estimate the number of clusters, providing a probabilistic guarantee against overestimation. Continuing with the example in Figure 1, we apply our proposed method to estimate the number of clusters, $\widehat{K}$, over $100$ runs of our randomized algorithm; panel (d) displays the empirical histogram of $\widehat{K}$, which has a strong mode at $2$, the true number of clusters.
@@ 1.1 Related work and contributions
Determining whether two clusters identified by a clustering algorithm are genuinely distinct, when the same data are used both to discover and to validate these clusters, is an instance of double dipping. Selective inferential methods address this by explicitly accounting for the dependence introduced by reusing the same data twice. Recent work by Gao et al. [2024] and Yun and Barber [2023] develop valid post-hoc inference on results obtained from agglomerative clustering. Beyond hierarchical clustering, Chen and Witten [2023] develop a selective inferential approach to validate clusters identified by the K-means algorithm. By conditioning on the output of the clustering algorithm, these approaches yield valid p-values at a prespecified dendrogram level, controlling the Type I error rate. Zhang et al. [2019] address a similar problem using sample splitting, which conditions on more information than necessary when compared with the aforementioned approaches. Other work, such as Liu et al. [2008], Maitra et al. [2012], Suzuki and Shimodaira [2006], test the global null hypothesis that no clustering structure exists in the data, which is fundamentally different in spirit from the hypotheses considered in the selective inference approaches.
To test the inherently data-dependent hypotheses at each merge in the dendrogram, we condition on the clustering outcome, following the conditional framework in existing selective inference methodologies for clustering. However, by introducing external randomization, our testing method differs considerably from previous approaches and offers several advantages, as outlined below.
@@
    Our approach produces substantially more powerful p-values than the non-randomized methods when testing mean differences between a pair of clusters at a fixed dendrogram height. This gain in power comes at little cost to the clustering quality when only a small amount of randomization is introduced, as seen in Figure 1, for example. Gains in inferential power due to external randomization have been well-documented in supervised learning contexts. Take, for example, data-carving methods in Fithian et al. [2014], Panigrahi [2023], which select models using a random subsample like sample-splitting, but use the full data for inference. Approaches using other types of randomization schemes are described in Tian and Taylor [2018], Panigrahi and Taylor [2023], Panigrahi et al. [2024], Kivaranovic and Leeb [2024], Huang et al. [2025], among others.
@@
    Furthermore, unlike randomized methods based on data-splitting, like the UV decomposition, fission, or thinning [Rasines and Young, 2022, Leiner et al., 2025, Dharamshi et al., 2025], that base inference exclusively on holdout data, our approach reuses the data employed in clustering. In particular, with a small amount of randomization, our approach produces only a slight perturbation of the clustering results on the original data while still using the full dataset for testing.
@@
    The p-values obtained from our randomized approach can be used in conjunction with any hierarchical clustering linkage to produce valid inference, without the need for: (i) case-by-case derivations of the conditioning set or (ii) potentially expensive Monte Carlo or resampling-based approximations to compute the p-values. For example, in complete-linkage hierarchical clustering (which was used in Figure 1), the tests in Gao et al. [2024], Yun and Barber [2023] rely on Monte Carlo approximations to compute p-values, whereas our method does not resort to such sampling-based approximations.
@@
    Adding to the flexibility of our approach, the proposed randomization scheme can be readily applied to obtain adjusted distributions of test statistics for different types of tests, while accounting for the data-dependent nature of clustering outputs. Examples include tests based on the adjusted chi-squared test statistic in Gao et al. [2024] and the adjusted F test statistic in Yun and Barber [2023].
@@
  • In Section 2, we introduce the randomized hierarchical clustering algorithm (Algorithm 1), which performs clustering using a dissimilarity matrix that is updated at each merge.
@@
  • Simulation results in Section 5 demonstrate that: (i) our randomized approach yields valid p-values at each merge in the dendrogram, (ii) tests based on these p-values achieve higher power than existing selective inference methods at a fixed dendrogram height, without sacrificing clustering performance when only a small amount of randomization is used, and (iii) the proposed $\alpha$-spending procedure to estimate the dendrogram cut height yields a powerful clustering method that outperforms the standard clustering paired with the popular gap statistic [Tibshirani et al., 2001]. In a real data analysis, we illustrate how the number of clusters estimated using our method can be used to assess the stability of clustering results across multiple runs of randomized Algorithm 1.
@@ 2 Randomized hierarchical clustering
In traditional agglomerative hierarchical clustering, we begin with each observation in its own cluster and at each step we find the pair of clusters that are closest and merge them. To evaluate each candidate merge $M=(C_{1},C_{2})$, we use a dissimilarity between the clusters $C_{1}$ and $C_{2}$ which in this context is called a linkage function, $d(M;X)$, where $X\in\mathbb{R}^{n\times p}$ is the data matrix. In particular, at step $t$ one chooses a merge
$M^{*(t)}\in\arg\min_{M\in\mathcal{M}^{*(t)}}d(M;X),$ (1)
where $\mathcal{M}^{*(t)}$ is the allowable set of merges at step $t$, consisting of all pairs of clusters existing at that step. In practice, there are two choices built into the function $d(M;X)$: (i) how to form a $n\times n$ dissimilarity matrix $D^{(1)}(X)$ between the rows of $X$ (e.g., Euclidean distance) and (ii) how to go from those base dissimilarities to the dissimilarity between a pair of clusters (e.g., taking the minimum/maximum/average dissimilarity between pairs of points in different clusters). Depending on these choices, ties may arise in (1), in which case there can be multiple merge sequences $\overline{M}^{*(t)}:=(M^{*(1)},\ldots,M^{*(t)})$ that are equally valid.
@@ 2.1 Proposed randomization scheme
Our proposed algorithm relaxes the greedy nature of traditional hierarchical clustering by randomizing the merge selection. Instead of selecting a minimizing merge $M^{*(t)}$ at each step, we sample a random merge $M^{(t)}$ from $\mathcal{M}^{(t)}$, the current set of allowable merges. Notationally, we remove stars to distinguish our random sequence from the greedy algorithm’s sequence. Furthermore, let $\overline{M}^{(t-1)}:=(M^{(1)},\ldots,M^{(t-1)})$ denote the random sequence of previous merges until step t1t-1. Then, note that $\mathcal{M}^{(t)}$, the set of allowable merges at step $t$, is a deterministic function of the past merges, i.e., $\mathcal{M}^{(t)}:=\mathcal{M}(\overline{M}^{(t-1)})$.
We are now ready to describe how we sample the random merge at step $t$. Conditional on $\overline{M}^{(t-1)}$ and $X$ (i.e, the sequence of previous merges and the data), $M^{(t)}$ is sampled according to the probability mass function
$\small\mathbb{P}\left(M^{(t)}=M|\overline{M}^{(t-1)}=\overline{M}_{o}^{(t-1)},X=X_{o}\right)=p^{(t)}(M;\mathcal{M}_{o}^{(t)},X_{o})=\dfrac{\exp\left(-d(M;X_{o})/\tau(\mathcal{M}_{o}^{(t)},X_{o})\right)}{\displaystyle\sum_{M^{\prime}\in\mathcal{M}_{o}^{(t)}}\exp\left(-d(M^{\prime};X_{o})/\tau(\mathcal{M}_{o}^{(t)},X_{o})\right)},$ (2)
where $X_{o}$ is the observed data matrix, $\tau$ is a user-specified function controlling the amount of external randomization in the merge selection, and $\mathcal{M}^{(t)}_{o}=\mathcal{M}(\overline{M}_{o}^{(t-1)})$ is the observed set of currently allowable merges.
In brief, we have replaced the $\arg\min$ over the vector $D^{(t)}(X):=\{d(M;X):M\in\mathcal{M}^{(t)}\}$ by sampling from the softmax of $D^{(t)}(X)/\tau(\mathcal{M}^{(t)},X)$. The randomized merging scheme introduced above closely resembles the exponential mechanism by McSherry and Talwar [2007], which is used in differential privacy to randomly sample an item from a discrete set while providing privacy guarantees. Recent work by Zhang et al. [2024] uses a similar type of sampling probabilities as weights in a weighted average-type procedure to obtain stability, which in turn is used to establish the asymptotic normality of cross-validated risks. A key distinction of our randomized approach from the broader literature on differential privacy and algorithmic stability is that we neither require nor seek to achieve any form of privacy or stability in order to construct valid testing procedures, as described later in Section 3.
@@ 2.2 Amount of randomization
At each step of the randomized clustering algorithm, one must determine the amount of randomization used to sample the merge. We define the user-specified function controlling the amount of external randomization as
$\tau(\mathcal{M}^{(t)},X):=\tau^{*}\cdot\frac{1}{|\mathcal{M}^{(t)}|}\sum_{M^{\prime}\in\mathcal{M}^{(t)}}d(M^{\prime};X),$ (3)
i.e., at step $t$, we set the amount of randomization to be a constant fraction $\tau^{*}$ of the average dissimilarity among the potential merges at that step. This data-driven specification of $\tau$ ensures that the clustering structure produced by our randomized algorithm remains invariant under rescaling of the data, a property also exhibited by the deterministic version.
The smaller the value of $\tau$, the lower the amount of randomization in our clustering algorithm. Our next result, Proposition 1, shows that as $\tau^{*}\to 0$, the random sequence of merges produced by our randomized clustering algorithm, with probability approaching $1$, coincides with the sequence generated by the traditional algorithm under the same linkage criterion. Corroborating this result, our numerical experiments in Section 5.1 show that small values of $\tau^{*}$ lead to clustering results that closely resemble the clustering output on the full data in terms of standard clustering evaluation metrics.
@@
Let $\Omega_{o}^{*(t)}$ be the set of possible merge sequences from running traditional hierarchical clustering (1) for $t$ steps on a fixed data matrix $X_{o}$, i.e.
$\Omega_{o}^{*(t)}=\left\{(\overline{M}^{*(t-1)}_{o},M^{*(t)}):\overline{M}^{*(t-1)}_{o}\in\Omega_{o}^{*(t-1)},~M^{*(t)}\in\operatorname*{arg\,min}_{M\in\mathcal{M}(\overline{M}^{*(t-1)}_{o})}d(M;X_{o})\right\}.$ (4)
Then, the randomized algorithm’s random merge sequence $\overline{M}^{(t)}$ when presented with the same data satisfies the following:
$\lim_{\tau^{*}\rightarrow 0}\mathbb{P}(\overline{M}^{(t)}\in\Omega^{*(t)}_{o}|X=X_{o})=1.$
[…]
Algorithm 1 presents the proposed randomized hierarchical clustering method, producing $K$ clusters by performing nKn-K merges. In the special case where $K=1$, the algorithm proceeds until all observations are merged into a single cluster.
@@
1:Input: Pairwise dissimilarity matrix $D^{(1)}(X)\in\mathbb{R}^{n\times n}$, number of clusters $K$, linkage function $d(\cdot;X)$, parameter controlling the amount of randomization $\tau^{*}$
2:Initialization: Each data point as its own singleton cluster: $\mathcal{C}^{(1)}=\left\{\{1\},\{2\},\dots,\{n\}\right\}$
@@
4:  Determine the set of allowable merges $\mathcal{M}^{(t)}$ from the clustering $\mathcal{C}^{(t)}$. For each pair of cluster $M\in\mathcal{M}^{(t)}$, compute $d(M;X)$ using linkage $d(\cdot;X)$.
5:  Set $\tau(\mathcal{M}^{(t)},X):=\tau^{*}\cdot\dfrac{1}{|\mathcal{M}^{(t)}|}\sum_{M\in\mathcal{M}^{(t)}}d(M;X)$
6:  Sample the winning merge $M^{(t)}=\left({C}^{(t)}_{1},{C}^{(t)}_{2}\right)$ from probability mass function (2).
7:  Merge the winning pair of clusters and update the collection of clusters:
$\mathcal{C}^{(t+1)}\leftarrow\mathcal{C}^{(t)}\setminus\{{C}^{(t)}_{1},{C}^{(t)}_{2}\}\cup\left\{{C}^{(t)}_{1}\cup{C}^{(t)}_{2}\right\}$
8:end for
9:Output: Cluster labels for each data point